OpenStudy (anonymous):

Show that if n and k are integers with 1 ≤ k ≤ n, then c(n,k)<=n^k/(2^(k-1)

OpenStudy (amistre64):

sounds inductive ...

OpenStudy (amistre64):

you know the formula for C(n,k) right?

OpenStudy (anonymous):

yes, I do

OpenStudy (anonymous):

I will give it a try by induction.Thanks though

OpenStudy (amistre64):

not sure if an algebra would be sufficient ... maybe

OpenStudy (anonymous):

I actually thought it requires a combinitorial proof.

OpenStudy (amistre64):

im not proficient enough to even recall what a combinatorial proof entails :) \[\frac{n!}{k!~(n-k)!}\le\frac{2n^k}{2^k}\]